OpenStudy (sanjanap):

An airbag was determined to have a maximum volume of 4.5 liters. If the most effective inflation percentage was 95% of the maximum, how much of each reactant below is needed at STP? Acetic acid: 0.250 molar. NaHCO3 + CH3COOH -->

OpenStudy (sanjanap):

@sleepyjess

OpenStudy (sanjanap):

@zepdrix

OpenStudy (sanjanap):

@dan815

OpenStudy (cuanchi):

NaHCO3 + CH3COOH= NaCH3COO + H20 + CO2 4.5 L x (95/100) x (1 mol CO2/ 22.4L) x (84.007g/1 mol NaHCO3) = ??? g of NaHCO3 4.5 L x (95/100) x (1 mol CO2/ 22.4L) x (60.05 g/ 1 mol CH3COOH) x (1L/0.250 molar) = ??? L of CH3COOH